OpenStudy (anonymous):

A tower casts a 450ft shadow at the same time a 4ft child casts a 6 ft shadow. What is the height of the tower

OpenStudy (anonymous):

\[\tan\theta = \frac{4}{6}\] \[\tan\theta = \frac{h}{450}\] \[\frac{4}{6}*450=h\]